# Differential Test Performance and Peer Effects

### v1

@@ 1 Introduction
We develop a framework to analyze, identify and estimate the effect of peer groups on performance measures for individuals who are allocated to groups. A leading example are test scores of students who are allocated to classrooms. However, the theory could equally be applied to outcome measures of members of sports teams, teams of workers or other groups. Identification is based on the availability of multiple performance measures of comparable quality or information content. We allow team formation to be endogenous but also cover the case of randomly selected groups. For expositional purposes we refer to groups as classrooms and individuals selected into groups as students with the understanding that our procedure applies to a broader set of scenarios. For the statistical implementation we impose linear functional form restrictions. This allows us to identify the marginal effect of observed and unobserved average peer characteristics.
@@
Our analysis centers around the idea of measuring the performance of individuals in closely related tasks and within a short period of time. An example are aptitude tests administered at the end of the school year and in related areas such as reading, writing and word comprehension. By considering the individual quasi-difference in scores we are able to eliminate unobserved ability and unobserved group effects. Our key identifying assumption postulates that different tests measure similar skills and that variation in differential scores is, apart from variation induced by observed covariates, due to cross-sectional variation in ability or effort of individuals that is not related to prior performance and other systematic and possibly unobserved factors correlated with group formation such as teacher quality or parental support. Our theory builds on the concept of latent performance, ability or effort measures defined in the absence of group interaction.
@@
The first result of our paper shows that our key identifying restriction implies orthogonality conditions that are at the core of our identification strategy. These restrictions are obtained without any assumptions about random group selection or independence of individual characteristics in the population of individuals. Combined with the assumption of linear latent outcomes and mean group peer effects this leads to an empirical model that falls within the class of linear peer effects models that have been prominently studied in the literature, see for example Manski (1993); Calvó-Armengol, Patacchini, and Zenou (2009); Blume et al. (2015) and Angrist (2014) for a critique of these models. We expand on this literature by explicitly accounting for endogenous peer group selection and individual heterogeneity in unobserved test taking ability. Most of the empirical literature uses observable characteristics as well as proxies for unobserved ability to measure the quality of peer groups. Our baseline model is formulated for unobserved, at least to the analyst, peer characteristics. We use multiple performance measures for similar skills to difference out common test taking ability. We allow group selection to depend on unobserved as well as observed characteristics. Recent contributions to the econometrics literature accounting for endogenous group and network formation include Goldsmith-Pinkham and Imbens (2013), Hsieh and Lee (2016) and Griffith (2022) who use an explicit network formation model, Qu and Lee (2015), Johnsson and Moon (2019), Auerbach (2022) who use control function approaches and Kuersteiner and Prucha (2020) who use an instrumental variables approach. [^1] The literature on peer effects is part of a larger literature on network effects. It is well recognized that the adjacency matrix frequently used in modeling peer effect is a special case of the weight matrices used in a class of models introduced by Cliff and Ord (1973, 1981), which were originally intended for modeling spatial network effects; see Anselin (2010) for a review of important contribution of this literature on identification and estimation of network effects, and Kuersteiner and Prucha (2020) for a recent contribution that connects those strands of network literature. .
In our setting, reduced form regression analysis that focuses on exogenous or contextual peer effects leads to biased estimates. We propose instrumental variables estimators to overcome these problems. By exploiting moment restrictions that are implied by our identifying assumptions we are proposing generalized method of moments (GMM) estimators that identify both exogenous as well as endogenous peer effects. Formally, our moments based estimator is a panel type estimator based on quasi-differences between test performance measures. An important difference between our approach and typical panel settings is the lack of a temporal dimension, and where the setup is geared towards the utilization of different test performance measures that are obtained in an essentially simultaneous fashion.
@@
The difficulties of identifying peer effects in models with transitive groups where all peers have the same effect, are well known since Manski (1993). Approaches to overcome identification challenges have focused on random group assignment, for example Sacerdote (2001); Duflo and Saez (2003); Carrell, Sacerdote, and West (2013); Sojourner (2013), exogenous variation in group size as in Lee (2007b) or Kuersteiner, Prucha, and Zeng (2023), instrumentation of peer effects using non-transitive group structures as in Bramoullé, Djebbari, and Fortin (2009), explicit group formation models in Goldsmith-Pinkham and Imbens (2013), variation in group specific heteroskedasticity in Graham (2008), and panel methods in Mas and Moretti (2009), Arcidiacono et al. (2012), Cornelissen, Dustmann, and Schönberg (2017), Miraldo, Propper, and Rose (2021) and Braun and Verdier (2023). The approach pursued in this paper builds on the dynamic panel methods for spatial and social networks developed in Kuersteiner and Prucha (2020).
We use a baseline model of peer effects that is similar to Arcidiacono et al. (2012) who consider peer effects working through unobserved individual characteristics. Arcidiacono et al. (2012) allow for individual specific effects as well as for endogenous group formation and use panel data. Technically, our estimator is different from theirs since we use a quasi-differencing approach to eliminate class and individual fixed effects while they use a non-linear fixed effects estimator. The more important difference lies in the assumptions and empirical implementation of the method we have in mind. While our estimator could be applied to more conventional panel data with performance measures observed at different points in time, we emphasize applications where measures of closely related skills are observed essentially simultaneously. For one, this framework alleviates problems with sample attrition which are well documented for Project STAR. More importantly, our core assumption on unobserved skill or effort centers around a lack of predictive power of additional test results for closely related test measures. For example, a student scoring well on a reading test is likely to score well on a word comprehension test given at the same time, and in this sense the word comprehension test does not reveal additional information about underlying skills or effort. An advantage of our framework is that its credibility can be assessed and influenced by a testing protocol that complies with its basic premise. This is in contrast to more detailed restrictions between various unobserved components as well as on the evolution of individual effects measuring the change in skills over time required in the framework of Arcidiacono et al. (2012). An additional advantage of our procedure is that it does not require estimating unobserved skill or effort measures of individuals.
@@
In our empirical work we use Project STAR data and focus on classrooms in kindergarten to the third grade. We consider four individual SAT scores, mathematics, reading, listening and word study skills. We document the high correlation between pairs of these scores, unconditionally as well as conditionally on a full set of controls, with correlation coefficients ranging from 0.5 to 0.9. The highest correlation is observed for the reading and word study skill scores. The correlation patterns are robust across the four grade levels we consider. The peer effects coefficient of our model is estimated precisely and is stable across the different grades with the exception of first grade where we observe a lower degree of peer effects. We find stronger peer effects for test score pairs that are less highly correlated, in line with our model assumptions where more closely related scores offer better control for unobserved confounding factors. Quantitatively, we estimate that an increase of average unobserved peer quality measured in SAT equivalent scores of 10 SAT points leads to an increase of 2 to 4 SAT points for an individual class mate. This estimate is based on the reading-word score pair. For other score pairs the effects range from a 4 to a 7 SAT point increase for the scores of individual classmates due to a change in average scores. These measurements are quite robust to the inclusion or omission of additional controls. When we measure exogenous peer effects, in other words the effects of average age, race and gender, as well as an indicator of economic background, we find mostly insignificant results on differential scores. This lack of significance is further evidence that our differencing strategy eliminates much of the systematic factors explaining test score variation. However, it does not mean that exogenous controls have no effect on the level of individual scores.
The paper is organized as follows. In Section 2 we develop the identification and estimation strategy. Section 2.1 defines performance measures and group selection, Section 2.2 introduces the education production function and peer effects, Section 2.3 discusses the key identifying restriction and Section 2.4 introduces the estimators we propose. Section 3 contains the empirical analysis of Project STAR data and Section 4 develops formal identification and consistency proofs as well as an asymptotic distribution theory. Section 5 contains conclusions. Proofs and tables are contained in the appendix.
@@ 2 Model
We assume that for each student $i=1,...,n\text{ }$we observe results for two different tests, or more generally that we observe two measures of academic achievement, say, $y_{it}$ where $t=1,2$ indexes the test. Our empirical work uses Project STAR data where these measures are scores of the Stanford Achievement Tests (SAT) chosen from math, word study skills, listening and reading tests given to kindergarten through third grade students. Our approach uses two outcome measures to control for unobserved and possibly correlated student characteristics as well as group level effects related to classrooms, teachers and schools.
@@ 2.1 Latent Performance and Group Selection
Selection into classrooms can be correlated with a latent baseline performance measure $\tilde{\zeta}_{i}^{*}$ that can also be thought of as prior test taking ability or latent performance, or effort prior to taking the observed tests. The purpose of introducing baseline test taking ability is to account for unobserved student characteristics that are invariant for the two tests. We account for observed personal student characteristics where $\tilde{v}_{i}^{p}$ and $\tilde{w}_{it}^{p}$ denote row vectors of observed test invariant and test varying characteristics. For notational convenience we collect the performance measures and characteristics for all $n$ students in the sample in the following vectors and matrices of row dimension $n$: $\tilde{\zeta}^{\ast}=(\tilde{\zeta}_{1}^{\ast},...,\tilde{\zeta}_{n}^{\ast})^{\prime}$, $\tilde{y}_{t}^{*}=\left(\tilde{y}_{1t}^{*},\ldots,\tilde{y}_{nt}^{*}\right)^{\prime}$ for $t=1,2$, $\tilde{v}^{p}=\left(\tilde{v}_{1}^{p\prime},\ldots,\tilde{v}_{n}^{p\prime}\right)^{\prime}$, $\tilde{w}_{t}^{p}=\left(\tilde{w}_{1t}^{p\prime},\ldots,\tilde{w}_{nt}^{p\prime}\right)^{\prime}$ for $t=1,2$, and $\tilde{X}^{p}=\left(\tilde{v}^{p},\tilde{w}_{1}^{p},\tilde{w}_{2}^{p}\right)$. The $i$-th row of $\tilde{X}^{p}$ then contains the covariates of student $i$ in the original sample. Selection into classrooms can, in addition to depending on $\tilde{\zeta}^{*}$ and $\tilde{X^{p}}$, depend on observed class level characteristics as well as on a vector $\alpha=\left(\alpha_{1},...,\alpha_{C}\right)^{\prime}$ of unobserved class level characteristics. We do not impose any restrictions on the cross-sectional dependence in $\tilde{\zeta}^{\ast}$, $\tilde{X}^{p}$, $\alpha$ or observed class level characteristics. Latent baseline performance of student $i$ may be correlated with latent performance of student $j$ for a variety of reasons including similar educational, socioeconomic, cultural, religious or geographic backgrounds that may or may not be known to the analyst.
@@
We now describe the selection of students to classrooms. This assignment process results in a reordering of the students characterized by a one-to-one mapping of the original index set to the new index set. Let $\tilde{y}_{t}^{*}=\left(\tilde{y}_{1t}^{*},...,\tilde{y}_{nt}^{*}\right)^{\prime}$ denote the vector of unobserved ability of students corresponding to the the original sample or population, let $S_{c}$ be the $n_{c}\times n\text{ }$ selection matrix that allocates students to classroom $c$, and let $S=\left[S_{1}^{\prime},...,S_{C}^{\prime}\right]^{\prime}$ be the selector matrix for all students. The set of all possible classroom allocations is denoted by $\mathscr{S}$. A random assignment mechanism can be thought of as selecting one element of $\mathscr{S}$ at random. Corresponding to $S$ the vector $y_{t}^{*}=\left(y_{1t}^{*},\ldots,y_{nt}^{*}\right)^{\prime}$ representing the sample or population of unobserved ability for test $t$ of students ordered by classrooms is given by
@@
Similarly, we arrange unobserved and observed student characteristics by classroom through the transformation $\zeta^{*}=\left(\zeta_{1}^{*},\ldots,\zeta_{n}^{*}\right)^{\prime}=S\tilde{\zeta}^{\ast}$, $v^{p}=\left(v_{1}^{p\prime},\ldots,v_{n}^{p\prime}\right)^{\prime}=S\tilde{v}^{p}$, $w_{t}^{p}=\left(w_{1t}^{p\prime},\ldots,w_{nt}^{p\prime}\right)^{\prime}=S\tilde{w}_{t}^{p}$ for $t=1,2$, and $X^{p}=\left(x_{1}^{p\prime},\ldots,x_{n}^{p\prime}\right)^{\prime}=S\tilde{X}^{p}=\left(v^{p},w_{1}^{p},w_{2}^{p}\right)$. Thus, the row vector $x_{i}^{p}$ now denotes observable personal characteristics for student $i$ in the sample ordered by classrooms.
We allow for $S$, $\tilde{\zeta}^{\ast}$, $\tilde{y}_{t}^{*}$, for $t=1,2$, and $\alpha$ to be mutually correlated and to be correlated with $\tilde{X}^{p}$ as well as with observed class level characteristics. This assumption includes scenarios where $S$ is selected completely at random, scenarios where $S$ is selected based on student and class level characteristics, as well as scenarios where students respond to the allocation $S$. Our assumptions, formally spelled out in Section 2.3, require that selection happens before testing which is a mild restriction that should be satisfied in most scenarios we have in mind. We do not otherwise specify or restrict the mechanism $S$ that selects students into classrooms. We also do not require that $S$ is observed. This is relevant in situations where we do not have information identifying individual students such as through their names, social security numbers or residential address. On the other hand we do require that we know who, among all test takers, is allocated to the same classroom. For project STAR data, this information is available for the majority of students in grades K through three through a class identifier [^2] The identifier is a teacher ID that is unique for each classroom., but not for students in higher grades.

### v2

@@ 1 Introduction
We develop a framework to analyze, identify and estimate the effect of peer groups on performance measures for individuals who are allocated to groups. Our analysis centers around the idea of measuring the performance of individuals in closely related tasks and within a short period of time. An example are aptitude tests administered at the end of the school year and in related areas such as reading, writing and word comprehension. By considering the individual quasi-difference in scores we are able to eliminate unobserved ability and unobserved group effects. Our key identifying assumption postulates that different tests measure similar skills and that variation in differential scores is, apart from variation induced by observed covariates, due to idiosyncratic cross-sectional variation in ability or effort of individuals. That variation is not related to prior performance and other systematic and possibly unobserved factors correlated with group formation such as teacher quality or parental support.
A leading example where our method can be applied are test scores of students who are allocated to classrooms. However, the theory could equally be applied to outcome measures of members of sports teams, teams of workers or other groups. Identification is based on the availability of multiple performance measures of comparable quality or information content. We allow team formation to be endogenous but also cover the case of randomly selected groups. For expositional purposes we refer to groups as classrooms and individuals selected into groups as students with the understanding that our procedure applies to a broader set of scenarios.
@@
The first result of our paper shows that our key identifying restriction implies orthogonality conditions that are at the core of our identification strategy. These restrictions are obtained without any assumptions about random group selection or independence of individual characteristics in the population of individuals. Combined with the assumption of linear latent outcomes and mean group peer effects this leads to an empirical model that falls within the class of linear peer effects models that have been prominently studied in the literature, see for example Manski (1993); Calvó-Armengol, Patacchini, and Zenou (2009); Blume et al. (2015) and Angrist (2014) for a critique of these models. We expand on this literature by explicitly accounting for endogenous peer group selection and individual heterogeneity in unobserved test taking ability. Most of the empirical literature uses observable characteristics as well as proxies for unobserved ability to measure the quality of peer groups. Our baseline model is formulated for unobserved, at least to the analyst, peer characteristics. We use multiple performance measures for similar skills to difference out common test taking ability. We allow group selection to depend on unobserved as well as observed characteristics. Recent contributions to the econometrics literature accounting for endogenous group and network formation include Goldsmith-Pinkham and Imbens (2013), Hsieh and Lee (2016) and Griffith (2022) who use an explicit network formation model, Qu and Lee (2015), Johnsson and Moon (2019), Auerbach (2022) who use control function approaches and Kuersteiner and Prucha (2020) who use an instrumental variables approach. [^1] The literature on peer effects is part of a larger literature on network effects. It is well recognized that the adjacency matrix frequently used in modeling peer effect is a special case of the weight matrices used in a class of models introduced by Cliff and Ord (1973, 1981), which were originally intended for modeling spatial network effects; see Anselin (2010) and Lee and Yu (2015) for a review of important contribution of this literature on identification and estimation of network effects from cross sectional and panel data, and Kuersteiner and Prucha (2020) for a recent contribution that connects those strands of network literature.
@@
Our estimators are related to methods that use panel data to identify peer effects, see Mas and Moretti (2009), Arcidiacono et al. (2012), Cornelissen, Dustmann, and Schönberg (2017), Miraldo, Propper, and Rose (2021) and Braun and Verdier (2023). In contrast with panel methods, we emphasize applications with cross-sectional data where measures of closely related skills are observed essentially simultaneously in time. For one, this framework alleviates problems with sample attrition which are well documented for Project STAR. More importantly, within this framework the source of identifying variation is due to remaining randomness, after conditioning on ability, in the responses of an individual to test questions rather than due to random selection of students into classrooms. An advantage of our framework is that its credibility can be assessed and influenced by a testing protocol that complies with its basic premise.
@@
In our empirical work we use Project STAR data and focus on students in kindergarten to the third grade. We use individual SAT scores in reading and word study skills as our outcome measures. We document the high correlation between the two scores, unconditionally as well as conditionally on a full set of controls. The correlation coefficients are about 0.9 for all four grades The peer effects coefficient of our model is estimated precisely and is stable across the different grades except for the first grade where we observe a lower degree of peer effects. Quantitatively, we estimate that an increase of average unobserved peer quality measured in SAT equivalent scores of 10 SAT points leads to an increase of 2 to 4 SAT points for an individual classmate. These measurements are quite robust to the inclusion or omission of additional controls. When we measure exogenous peer effects, in other words the effects of average age, race and gender, as well as an indicator of economic background, we find mostly insignificant results on differential scores. This lack of significance is further evidence that our differencing strategy eliminates much of the systematic factors explaining test score variation. However, it does not mean that exogenous controls have no effect on the level of individual scores.
The paper is organized as follows. In Section 2 we develop the identification strategy for a model without covariates. Section 2.1 defines performance measures and group selection, Section 2.2 introduces the education production function and peer effects and Section 2.3 discusses the key identifying restriction that defines closely related skills. Section 3 extends the specification to include covariates and introduces the estimators we propose. Section 4 contains the empirical analysis of peer effects using Project STAR data. We report a battery of specification checks in Section 4.2. Our main empirical findings are reported in Section 4.3. Section 5 develops formal identification proofs. Section 6 contains conclusions. Proofs and tables are contained in the appendix. An online appendix contains additional lemmas used in the proofs. In the online appendix we also give results that formally establish the consistency and asymptotic normality of the considered estimators, as well as some additional empirical results.
@@ 2 Model
We assume that for each student $i=1,...,n\text{ }$we observe results for two different tests, or more generally that we observe two measures of academic achievement, say, $y_{it}$ where $t=1,2$ indexes the test. Our empirical work uses Project STAR data where these measures are scores of the Stanford Achievement Tests (SAT) in word study skills and reading tests given to kindergarten through third grade students. Our approach uses two outcome measures to control for unobserved and possibly correlated student characteristics as well as group level effects related to classrooms, teachers and schools.
@@ 2.1 Latent Performance and Group Selection
Selection into classrooms can be correlated with a latent baseline performance measure $\tilde{\zeta}_{i}^{*}$ that can also be thought of individual test taking ability or latent performance, or effort prior to taking the observed tests. The purpose of introducing baseline test taking ability is to account for unobserved student characteristics that are invariant for the two tests. Our key identifying assumption of the two tests measuring essentially identical skills detailed in Section 2.3 implies that the ability $\tilde{y}_{it}^{*}$ of taking test $t$ can be decomposed into test invariant factors that include baseline ability $\tilde{\zeta}_{i}^{*}$ and idiosyncratic errors that are uncorrelated across students and tests. Our method also accounts for observable student and classroom specific characteristics. We omit the observed covariates for now to simplify the discussion of our identification strategy. We discuss a more complete model that includes covariates in Section 3. For notational convenience we collect the latent baseline performance measures for all $n$ students in the sample in the following $n\times 1$ vector: $\tilde{\zeta}^{\ast}=\left(\tilde{\zeta}_{1}^{\ast},...,\tilde{\zeta}_{n}^{\ast}\right)^{\prime}$. Selection into classrooms can, in addition to depending on $\tilde{\zeta}^{*}$ and observed individual characteristics, depend on observed class level characteristics as well as on a vector $\alpha=\left(\alpha_{1},...,\alpha_{C}\right)^{\prime}$ of unobserved class level characteristics. We do not impose any restrictions on the cross-sectional dependence in $\tilde{\zeta}^{\ast}$, $\alpha$ or observed individual and class level characteristics. Latent baseline performance of student $i$ may be correlated with latent performance of student $j$ for a variety of reasons including similar educational, socioeconomic, cultural, religious or geographic backgrounds that may or may not be known to the analyst.
@@
We now describe the selection of students into classrooms. This assignment process results in a reordering of the students characterized by a one-to-one mapping of the original index set to the new index set. Let $\tilde{y}_{t}^{*}=\left(\tilde{y}_{1t}^{*},...,\tilde{y}_{nt}^{*}\right)^{\prime}$ denote the vector of unobserved test-taking ability of students taking test $t$. The vector $\tilde{y}_{t}^{*}$ is indexed corresponding to the the original sample or population. Let $S_{c}$ be the $n_{c}\times n\text{ }$ selection matrix that allocates students to classroom $c$, and let $S=\left[S_{1}^{\prime},...,S_{C}^{\prime}\right]^{\prime}$ be the selector matrix for all students. The set of all possible classroom allocations is denoted by $\mathscr{S}$. A random assignment mechanism can be thought of as selecting one element of $\mathscr{S}$ at random. Corresponding to $S$ the vector $y_{t}^{*}=\left(y_{1t}^{*},\ldots,y_{nt}^{*}\right)^{\prime}$ representing the sample or population of unobserved ability for test $t$ of students ordered by classrooms is given by
@@
The latent performance vector $\zeta^{*}$ is defined analogously.
We allow for $S$, $\tilde{\zeta}^{\ast}$, $\tilde{y}_{t}^{*}$, for $t=1,2$, and $\alpha$ to be mutually correlated and also to be correlated with observed individual and class level characteristics for the generalized model presented in Section 3. This assumption includes scenarios where $S$ is selected completely at random, scenarios where $S$ is selected based on student and class level characteristics, as well as scenarios where students respond to the allocation $S$. Our assumptions, formally spelled out in Section 2.3, require that selection happens before testing which is a mild restriction that should be satisfied in most scenarios we have in mind. We do not otherwise specify or restrict the mechanism $S$ that selects students into classrooms. We also do not require that $S$ is observed. This is relevant in situations where we do not have information identifying individual students such as through their names, social security numbers or residential address. On the other hand we do require that we know who, among all test takers, is allocated to the same classroom. For project STAR data, this information is available for the majority of students in grades K through three through a class identifier [^2] The identifier is a teacher ID that is unique for each classroom., but not for students in higher grades.